OpenStudy (smurfy14):

Solve using algebra. 6x^2-13x+6≥0

OpenStudy (anonymous):

\[(3x-2)(2x-3)\ge\] \[x \ge2/3, x \ge3/2\]

OpenStudy (anonymous):

start with \[(2x-3)(3x-2)\geq 0\] then see that the zeros are at \[\frac{2}{3}\] and \[\frac{3}{2}\] since you have a quadrant with positive leading coefficient, it is a parabola that opens up so it will be positive outside the roots and negative between them. you want positive, so answer is \[(-\infty, \frac{2}{3}]\cup (\frac{3}{2},\infty)\]

OpenStudy (anonymous):

*quadratic, not "quadrant"

OpenStudy (smurfy14):

i dont get why it is negative infinity to 2/3???

OpenStudy (anonymous):

because it is every number except those that we found

OpenStudy (anonymous):

\[y=6x^2-13x+6\] looks like this 

OpenStudy (anonymous):

it is negative (below the x axis) between the zeros and positive (above the x axis) outside the zeros

OpenStudy (smurfy14):

ya but if not looking at a graph, how would you know?

OpenStudy (anonymous):

you don't actually have to draw the picture, that is just my explanation. a quadratic with positive leading coefficient will be negative between the zeros and positive elsewhere

OpenStudy (anonymous):

all quadratics look the same. you will not get a different picture. it will either open facing up or facing down depending on the leading coefficient

OpenStudy (smurfy14):

ohhhh ok i get it now i see. thanks satellite :)

OpenStudy (anonymous):

yw